# Hyperparameter Optimization for Randomized Algorithms: A Case Study for Random Features

### v2

@@ 4.1.2 Sobol G Function
The Sobol G function (G. E. B. Archer and Sobol, 1997; Saltelli et al., 2010), defined by
@@
is another GSA test function that additionally allows for varying input dimension $d$ and for tuneable interaction between different inputs via the choice of coefficients $\{a_{i}\}_{i=1}^{d}$. Here we take $a_{i}=\frac{i-1}{2}\geq 0$, where small $i$ (and thus small $a_{i}$) imply larger first-order effects, and interactions are primarily present between these variables.
For this experiment, the results are only presented with a trained RF. The GP was accurate to $\mathsf{V}$-approx for dimensions $d=3,6$, and $10$, but ran out of memory for $d=20$. The untrained RF performed very poorly and was removed for plot readability. Over input dimensions $d=3,6,10$, and $20$, approximately $2000\cdot d$ Sobol points are used to produce empirical estimates of first-order Sobol indices (exact number determined by the GSA package). Of these, $250\cdot d$ (around 10-15%) were used for training the emulator. The posterior mean evaluation at all Sobol points was then used to calculate empirical Sobol indices.
Figure 2 shows the indices $\mathsf{V}_{i}$ and $\mathsf{TV}_{i}$ for $i=1,\dots d$, with one panel for each experiment. The emulation was repeated 30 times to produce the spread. Qualitatively, the RF is always able to capture the decaying sensitivity as $i$ increases, and quantitatively for $d=3,6$, and $10$, the RF spread well captures the analytic indices, while for $d=20$, only the $\mathsf{V}_{i}$ index is captured well. Most noticeably for the $\mathsf{TV}_{i}$ indices, there is a dimensional effect: as more insensitive dimensions are added, the RF emulation begins to overpredict the $\mathsf{TV}_{i}$ of the least sensitive dimensions and correspondingly overpredicts $\mathsf{TV}_{i}$ in the most sensitive dimensions.
Figure 3 provide views in the three most sensitive dimensions for just one of the RF realizations. The emulator has more of a challenge to capture the sharp transition at $x_{i}=0.5$ in the most sensitive dimensions as $d$ increases.
Figure 2: First order global sensitivity analysis of the Sobol G function with input dimensions $d=3,6,10,20$. Crosses ($\times$) denote the analytic Sobol indices, circles ($\circ$) denote the empirical indices calculated at $1600\cdot d$ points, error bars denote the $(0.05,0.95)$ percentile range of 30 random feature trials, trained on $250\cdot d$ points.
Figure 3: views of the Sobol G function plotted against the first three variables, for $d=3,6,10,20$ (top to bottom row). The red points denote the noisy observed data, and the blue represents the emulator prediction at all other sobol points.
@@ 4.2 Lorenz 63 Integrator
The next task is to learn an integrator, or $\mathbb{R}^{3}\to\mathbb{R}^{3}$ next-step map, of the Lorenz 63 system (Lorenz, 1963). The data is produced from observations of a trajectory of the full state, perturbed with additive Gaussian noise. There is growing use of the Lorenz system (without noise) in the recurrent network literature to validate the learning of deep or recurrent structure from full-state observations; here learning is demonstrated efficiently without deep networks. As the Lorenz 63 system gives rise to the archetypal chaotic dynamics, the measure of a successful integrator is an accurate representation of the chaotic attractor over long integration times, rather than fitting to a test trajectory.
@@
The classical parameter choices $(\sigma,\beta,\rho)=(10,\frac{8}{3},28)$ are chosen for the Lorenz ODE system to produce chaotic dynamics. The state is denoted by$(x(t),y(t),z(t))$ at each time $t$. The true dynamics were integrated with an Euler method and timestep of $0.01$. For the sample size, $500$ training data pairs were sourced randomly [^9] In experiments, sourcing data sequentially from a trajectory gives similar performance and introducing a spin-up time (e.g., learn with data from $T=[10,20]$) decreased performance slightly from $[0,20]$. The data pairs therefore comprise inputs $\{(x(t_{n}),y(t_{n}),z(t_{n})\}$ and (noisy) outputs $\{(x(t_{n+1}),y(t_{n+1}),z(t_{n+1}))+\eta\}$, where $\eta\sim\mathsf{N}(0,\Sigma)$. We emulate these maps with both scalar GPs (requiring decorrelation of the output space) and with vector-valued RFs (native output space).
For validation, we recursively apply the emulated integrators on time interval $[20,1020]$, plotting the trace and emulated attractor over $[20,120]$ and the marginal PDFs of the individual variables. The results for $\Sigma=10^{-4}I$ are shown in Figure 4, with the true integrated system in orange and the emulated system in blue. Row 1 shows the fit of $600$ untrained random features. Row 2 shows the fit of the GP emulator, and Row 3 shows the fit of the vector-valued RF emulation. Without hyperparameter learning, the integrator failed to reproduce the attractor, while both vector-valued RF and GP methods demonstrate stable integration for long times. The trace plot illustrates a forecast skill until $T=4$ for RF and $T=6$ for GP, and after these horizons the trajectories diverge as expected. Importantly, the emulated integrators continue to remain on the attractor, fitting the long-time PDFs closely. The approach is relatively robust to noise in the observations, as seen in the supplementary Figure 9, where the noise of the data is increased to $\Sigma=10^{-2}I$. Here the short term forecasts are still relatively good and the dynamics still explore the chaotic attractor. However, the long-time PDFs lose accuracy. Further increases to the noise eventually lead to a catastrophic failure, where emulators were observed to collapse to periodic orbits or fixed points.
@@
Figure 4: Learning a Lorenz 63 integrator from noisy data; emulated (blue) and truth (orange). Row 1 shows integration with the untrained RF, without hyperparameter learning. Row 2 shows integration with the GP emulator with 12 total hyperparameters learnt, and Row 3 shows integration with RF using a rank-4 nonseparable kernel with 31 total hyperparameters learnt. The $500$ data pairs, were subjected to observational noise with covariance $\Sigma=10^{-4}I$
@@
Figure 5: Comparison of true marginal empirical CDF (orange) with the CDF of 30 retrainings of the hyperparameters (blue) of the $\Sigma=10^{-4}I$.
@@
To inspect the robustness of the optimization methods: Figure 5 shows the difference between the true and emulated CDFs over 30 reruns of the random optimization algorithm. Parameter sweeps also show that the algorithm is robust to the data seeding, the number of features used for optimization and prediction, and to various modifiers of the optimization algorithm. Our results are in-line with efficient integrator learning approaches for this chaotic system, e.g., recurrent networks, reservoir computers, or vector auto-regression models (Shahi et al., 2022) that are often trained on $10^{3}$ to $10^{4}$ noiseless points, while feedforward or deep neural networks are frequently trained on more than $10^{4}$ noiseless points (Dudul, 2005; Fablet et al., 2018; Scher and Messori, 2019). Further details of the implementation of this experiment relating to EKI configuration and convergence are provided in Appendix A.4.
4.3 Accelerated Uncertainty Quantification for a Cloud Parametrization Scheme
@@
The final task is to investigate a parametric uncertainty quantification problem. Here, one has access to a black-box computational simulator (specifically a parameter-to-data map) $G$ that is dependent on several parameters (e.g., unknown physical constants) $\theta$, and the task is, given a noisy observation $y$, to estimate the posterior parameter distribution that produced this data. Common challenges in scientific realization of such problems are that (1) the simulator can only be run relatively few times (say, $<1000$), (2) the parameter-to-data map is likely non-differentiable or noisy, or both, and (3) the parametrizations are not directly observable. One can address these challenges by posing the underlying task as a classical Bayesian inverse problem and using gold-standard sampling methods such as Markov Chain Monte Carlo (MCMC) (Metropolis et al., 1953; Sherlock et al., 2010). MCMC algorithms draw samples from a density proportional to $\exp(-L(\theta,y))$, where the negative log-posterior $L$ is given by
$L(\theta,y)=\frac{1}{2}\|\Gamma^{-1/2}(y-G(\theta))\|^{2}+\frac{1}{2}\log\det\Gamma-\log P(\theta),$ (51)
@@
where $P(\theta)$ denotes the probability density function of the prior on $\theta$. Unfortunately, MCMC sampling algorithms may require $10^{3}\times$ more model evaluations than those available. Furthermore, non-differentiable and noisy forward maps may cause rough landscapes that are challenging to sample the posterior from.
@@
One approach to address the challenges of using MCMC is to replace the parameter-to-data (forward) map with a smooth and fast-evaluating surrogate (i.e., a statistical machine learning emulator) and sample using the surrogate instead. For example, when the forward map is random, previous work (Cleary et al., 2021; Dunbar et al., 2021) approximated samples of the forward map by a trained Gaussian process $G\sim\mathsf{GP}(G_{\mathrm{ML}},\Gamma_{\mathrm{ML}})$. Then, (51) can be replaced with
@@
$L(\theta,y)=\frac{1}{2}\|(\Gamma_{\mathrm{ML}}(\theta))^{-1/2}(y-G_{\mathrm{ML}}(\theta))\|^{2}+\frac{1}{2}\log\det\Gamma_{\mathrm{ML}}(\theta)-\log P(\theta)$ (52)
@@
Such an approximation and efficient selection of training points around the posterior mass is discussed at length in these studies; the algorithm is known as Calibrate-Emulate-Sample (CES) (Cleary et al., 2021). The statistical emulator of the forward map provides a natural smoothing property as it can directly predict the mean of the random process. In the present investigation, the goal is to compare the resulting approximate posterior distributions for a scientific inverse problem using GP emulators (GP-CES), and RF emulators (RF-CES).
@@
The example in this section is taken from the geophysics literature (Lopez-Gomez et al., 2022). At its core, the emulation task involves representing a map $G\colon\mathbb{R}^{5}\to\mathbb{R}^{50}$. To handle the multi-output problem with GPs, the output space is whitened and fifty independent $\mathbb{R}^{5}\to\mathbb{R}$ maps are trained. This approach depends entirely on the existence of this whitening approximation (which assumes a statistical structure on the inputs and outputs) and the training data sample size, while the new proposed approach using vector-valued RFs does not.
@@
In this example, the forward model is a one-dimensional vertical column model that represents turbulence and moist convection processes in the atmosphere, and it can represent cloud formation. The model is a parameterization of three-dimensional flow and is a type of eddy-diffusivity mass-flux (EDMF) model. It depends on several parameters governing eddy dissipation and diffusion due to turbulence, stability under convection, and the transport of tracers in (entrainment) and out (detrainment) of cloud updrafts. Both the model and experiment setup is detailed at length in (Lopez-Gomez et al., 2022). Data is taken to be time-averaged liquid water path (a scalar value measuring the amount of cloud water in the column) from a suite of $50$ three-dimensional cloud-resolving Large Eddy Simulations (LES); each LES simulation represents a different forcing scenario that mimics a range of sites and climates. The observational noise comes from the LES and uses artificial noise inflation to to reasonably capture the structural model error. EKI was used to find optimal parameter values, and in this companion experiment we use an emulator trained off of the EKI input-output samples within CES to approximate the full posterior distribution.
@@
Figure 6: Application of CES to the EDMF problem. Displayed are corner-plots showing slices through the joint distribution of five physically meaningful parameters calibrated to data. Green scatter: samples of the approximate posterior distribution obtained using RF-CES with no hyperparameter training; orange (blue) contours: the density of the posterior distribution obtained using GP-CES (RF-CES) after training. Both panels display the same experiment, but the right panel has the untrained emulator removed and axes rescaled.
@@
Figure 6 displays pairwise slices through the five-dimensional posterior distribution for the geophysical problem. Visualized are the final $10^{5}$ steps of MCMC using the likelihood calculated from (52). The green scatter shows posterior distribution samples resulting from an RF-CES strategy without hyperparameter training. The samples cover much of the the geophysical parameter prior support (see Table 3). In orange contours, the GP-CES algorithm is used. In blue contours the RF-CES algorithm is used. The left and right panels show the same experiment, just without the untrained samples. In the left panel, good qualitative agreement is shown across the support of the posterior from either approach and a significant improvement over the distribution from the untrained emulator. In the right panel, the GP-CES distribution can be seen to contain some fine-scale artifacts, while the RF distribution appears generally smoother. Such artifacts appear to be due to learning of small-lengthscales in the GP hyperparameter training algorithm rather than physically meaningful features in the posterior distribution. The GP-CES posterior was also slower to converge, and four times more samples of burn-in were taken in Figure 6 than in RF-CES. Likewise, pairwise correlation between several variables such as entrainment (entrainment_factor) and the stability coefficient (static_stab_coeff) are more clearly seen in the vector-valued RF setting. Further details of the implementation of this experiment relating to EKI configuration and convergence are provided in Appendix A.5.

### v4

@@ 4.1.2 Sobol G Function
The Sobol G function (Archer et al., 1997; Saltelli et al., 2010), defined by
@@
is another GSA test function that additionally allows for varying the input dimension $d$ and for tuneable interaction between different inputs via the choice of coefficients $\{a_{i}\}_{i=1}^{d}$. Here we take $a_{i}\coloneqq(i-1)/2\geq 0$, where small $i$ (and thus small $a_{i}$) imply larger first-order effects; interactions are primarily present between these variables.
For this experiment, the results are only presented with a tuned RF. We remark that a tuned GP was accurate in offline tests for dimensions $d=3,6$, and $10$, but timed out at $d=20$ dimensions. The untuned RF performed very poorly and was removed for plot readability. Over input dimensions $d=3,6,10$, and $20$, approximately $2000\cdot d$ Sobol points are used to produce empirical estimates of first-order Sobol indices. The exact number is determined by the GSA package GlobalSensitivityAnalysis.jl v1.2.0. Of these points, $250\cdot d$ (approximately $10\%$ to $15\%$) were used for tuning the emulator. The posterior mean evaluation at all Sobol points was then used to calculate empirical Sobol indices.
Figure 2: First order global sensitivity analysis of the Sobol G function with input dimensions $d=3,6,10,20$. The vertical axis represents the dimensionless values and the horizontal axis represents the index. Crosses ($\times$) denote the analytic Sobol indices, circles ($\circ$) denote the empirical indices calculated at $1600\cdot d$ points, error bars denote the $(0.05,0.95)$ percentile range of $30$ random feature trials with $250\cdot d$ training data points.
Figure 2 shows the indices $\mathsf{V}_{i}$ and $\mathsf{TV}_{i}$ for $i=1,\dots d$ with one panel for each experiment. The emulation was repeated 30 times to produce the spread. Qualitatively, the RF is always able to capture the decaying sensitivity as $i$ increases. Quantitatively, for $d=3,6$, and $10$, the RF spread well captures the analytic indices, while for $d=20$, only the $\mathsf{V}_{i}$ index is captured well. Most noticeably for the $\mathsf{TV}_{i}$ indices, there is a dimensional effect: as more insensitive dimensions are added, the RF emulation begins to overpredict the $\mathsf{TV}_{i}$ of the least sensitive dimensions and correspondingly overpredicts $\mathsf{TV}_{i}$ in the most sensitive dimensions. Figure 3 provide views in the three most sensitive dimensions for just one of the RF realizations. The emulator has more of a challenge to capture the sharp transition at $x_{i}=0.5$ in the most sensitive dimensions as $d$ increases. Appendix A.3 contains additional results pertaining to emulator performance as the input dimension $d$ and the number of features $M$ are varied.
Figure 3: Views of the Sobol G function plotted against the first three variables for $d=3$, $6$, $10$, and $20$ (from top to bottom row). The red points denote the noisy observed data, and the blue represents the tuned RF emulator prediction at all other Sobol points.
@@ 4.2 Lorenz 63 Integrator
The next task is to learn an integrator, or $\mathbb{R}^{3}\to\mathbb{R}^{3}$ next-step map, of the Lorenz 63 dynamical system (Lorenz, 1963). The data is produced from observations of a trajectory of the full state that is perturbed with additive Gaussian noise. There is growing use of the Lorenz system (without noise) in the recurrent network literature to validate the learning of deep or recurrent structure from full-state observations (Shahi et al., 2022; Dudul, 2005; Fablet et al., 2018; Scher and Messori, 2019); here, learning is demonstrated efficiently without deep networks. Since the Lorenz 63 system gives rise to the archetypal chaotic dynamics, the measure of a successful integrator is an accurate representation of the chaotic attractor over long integration times rather than pointwise accurate fits to a test trajectory.
The classical parameter choices $(\sigma,\rho,\beta)\coloneqq(10,28,8/3)$ are chosen for the Lorenz system
$\frac{dx}{dt}$ $=\sigma(y-x)\,,$ (51a)
$\frac{dy}{dt}$ $=x(\rho-z)-y\,,$ (51b)
$\frac{dz}{dt}$ $=xy-\beta z$ (51c)
@@
to produce chaotic dynamics. The state is denoted by $(x(t),y(t),z(t))$ at each time $t$. The true dynamics were integrated with an Euler method and time step of $10^{-2}$. For the sample size, $500$ training data pairs were sourced randomly [^9] In experiments, sourcing data sequentially from a trajectory gives similar performance and introducing a spin-up time (e.g., learn with data from time interval $[10,20]$) decreases performance slightly. from $[0,20]$. The data pairs therefore comprise inputs $\{(x(t_{n}),y(t_{n}),z(t_{n})\}$ and (noisy) outputs $\{(x(t_{n+1}),y(t_{n+1}),z(t_{n+1}))+\eta\}$, where $\eta\sim\mathsf{N}(0,\Sigma)$ and $\{t_{n}\}$ is the Euler time discretization. We emulate these maps with both scalar GPs (requiring decorrelation of the output space) and with vector-valued RFs (native output space).
@@
Figure 4: Learning a Lorenz 63 integrator from noisy data: emulated (blue) vs. truth (orange). The first column displays the time evolution of the three state variables $x(t)$, $y(t)$, and $z(t)$ (vertical axis) as a function of time $t$ (horizontal axis) on initial conditions not seen during training. The second column visualizes the marginal probability density functions for each state variable. Row 1 shows integration with the untuned RF without hyperparameter learning. Row 2 shows integration with the GP emulator with 12 total hyperparameters learned, and Row 3 shows integration with RF using a rank-$4$ nonseparable feature distribution with 31 total hyperparameters learned. The $500$ data pairs were subjected to observational noise with covariance $\Sigma=10^{-4}I$.
For validation, we recursively apply the emulated integrators on time interval $[20,1020]$, plotting the trace and emulated attractor over $[20,40]$ and the marginal probability density functions (PDFs) of the individual state variables. The results for $\Sigma=10^{-4}I$ are shown in Figure 4, with the true integrated system in orange and the emulated system in blue. Row 1 shows the fit of an untuned RF emulator with $M=600$ features. Row 2 shows the fit of the tuned GP emulator, and Row 3 shows the fit of the tuned vector-valued RF emulation. Without hyperparameter learning, the integrator failed to reproduce the attractor, while both the tuned vector-valued RF and GP methods demonstrate stable integration for long times. The trace plot illustrates a trajectory forecast skill until time $T=4$ for RF and for GP, and after these time horizons the trajectories diverge as expected. Importantly, the emulated integrators continue to remain on the attractor, fitting the long-time PDFs closely. The approach is relatively robust to noise in the observations, as seen in the supplementary Figure 9, where the noise of the data is increased to $\Sigma=10^{-2}I$. Here the short term forecasts are still relatively accurate and the dynamics still explore the chaotic attractor. However, the long-time PDFs lose accuracy. Further increases to the noise eventually lead to a catastrophic failure, where emulators were observed to collapse to periodic orbits or fixed points.
@@
Figure 5: Comparison of true marginal empirical CDF (orange) with the CDF of the RF emulator corresponding to $30$ re-tunings of the hyperparameters (blue) when performing the experiment displayed in Figure 4. The noise covariance is $\Sigma=10^{-4}I$.
@@
To inspect the robustness of the optimization methods, Figure 5 shows the difference between the true cumulative density function (CDF) and emulated CDF corresponding to the tuned RF emulator over 30 independent runs of the stochastic optimization algorithm. Parameter sweeps also show that the algorithm is robust to the data seeding, to the number of features used for optimization and prediction, and to various modifiers of the optimization algorithm. Our results are consistent with efficient integrator learning approaches for this chaotic system. These approaches include recurrent networks, reservoir computers, or vector auto-regression models (Shahi et al., 2022) that are often tuned and fit to $10^{3}$ to $10^{4}$ noiseless points. In contrast, feedforward or deep neural networks are frequently trained on more than $10^{4}$ noiseless points (Dudul, 2005; Fablet et al., 2018; Scher and Messori, 2019). Further details about the implementation of this experiment relating to EKI configuration and its convergence are provided in Appendix A.4. This appendix also includes supplementary results that explore the effect of feature distribution structure—such as rank and separability—on performance.
4.3 Uncertainty Quantification of a Cloud Model
@@
The final task is to investigate a parametric uncertainty quantification problem. Here, one has access to a black-box computational simulator $G$—specifically a parameter-to-data map—that is dependent on several parameters $\theta$, e.g., unknown physical constants. Given a noisy observation $y$, the task is to estimate the posterior parameter distribution that explains this measured data. Common challenges in the scientific realization of such problems are that (i) the simulator can only be run a relatively few times (e.g., fewer than $10^{3}$ times), (ii) the parameter-to-data map is likely non-differentiable or noisy, or both, and (iii) the parametrizations are not directly observable. One can address these challenges by posing the underlying task as a classical Bayesian inverse problem and using gold-standard sampling methods such as MCMC (Metropolis et al., 1953; Sherlock et al., 2010). Under a Gaussian noise model with covariance matrix $\Gamma$, [^10] This matrix $\Gamma$ is inherent to the physical inverse problem and should not be confused with the noise model for the EKI algorithm appearing in (45), which includes a matrix denoted by the same symbol. MCMC algorithms draw samples from a density over $\theta$ proportional to $\exp(-L(\theta,y))$, where the negative log-posterior is proportional to
$L(\theta,y)\coloneqq\frac{1}{2}\big{\lVert}\Gamma^{-1/2}\bigl{(}y-G(\theta)\bigr{)}\big{\rVert}^{2}+\frac{1}{2}\log\det\Gamma-\log P(\theta).$ (52)
@@
In (52), $P(\theta)$ denotes the probability density function of the prior on $\theta$. Unfortunately, MCMC sampling algorithms may require $\mathcal{O}(10^{3})$ times more model evaluations than those available. Furthermore, non-differentiable and noisy forward maps $G$ may cause rough landscapes that are challenging to sample the posterior from.
@@
One approach to address the challenges of using MCMC is to replace the parameter-to-data forward map with a smooth and fast-to-evaluate surrogate (e.g., a statistical machine learning emulator) and sample using the surrogate instead. For example, when the forward map is random, previous work (Cleary et al., 2021; Dunbar et al., 2021) approximated the forward map by a tuned GP posterior $G\sim\mathsf{GP}(G_{\mathrm{ML}},\Gamma_{\mathrm{ML}})$. Then, (52) can be replaced with
@@
$L(\theta,y)\coloneqq\frac{1}{2}\big{\lVert}\bigl{(}\Gamma_{\mathrm{ML}}(\theta)\bigr{)}^{-1/2}\bigl{(}y-G_{\mathrm{ML}}(\theta)\bigr{)}\big{\rVert}^{2}+\frac{1}{2}\log\det\bigl{(}\Gamma_{\mathrm{ML}}(\theta)\bigr{)}-\log P(\theta).$ (53)
@@
This approximation and the efficient selection of training points around the regions of high posterior mass are discussed at length in these studies; the algorithm is known as Calibrate-Emulate-Sample (CES) (Cleary et al., 2021). The statistical emulator of the forward map provides a natural smoothing property as it can directly predict the mean of the random process. In the present paper, the goal is to compare the resulting approximate posterior distributions for a scientific inverse problem using GP emulators (GP-CES) and RF emulators (RF-CES).
@@
The example in this section is taken from the geophysics literature (Lopez-Gomez et al., 2022). At its core, the emulation task involves representing a map $G\colon\mathbb{R}^{5}\to\mathbb{R}^{50}$. To handle the multi-output problem with GPs, the output space is whitened and $50$ independent $\mathbb{R}^{5}\to\mathbb{R}$ maps are tuned and fitted. The $50$ GPs can be viewed as a single vector-valued GP with a corresponding diagonal matrix-valued kernel. This approach depends entirely on the existence of the whitening approximation—which assumes a statistical structure on the inputs and outputs—and the training data sample size, while the new proposed approach using vector-valued RFs does not.
@@
In this scientific application, the forward map is a one-dimensional vertical column model that represents turbulence and moist convection processes in the atmosphere. It can also represent cloud formation. The model is a parametrization of three-dimensional flow and is a type of eddy-diffusivity mass-flux (EDMF) model. It depends on several parameters governing eddy dissipation and diffusion due to turbulence, stability under convection, and the transport of tracers in (entrainment) and out (detrainment) of cloud updrafts. Both the model and experiment setup is detailed at length in the work of Lopez-Gomez et al. (2022). The data is taken to be time-averaged liquid water path (a scalar value measuring the amount of cloud water in the column) from a suite of $50$ three-dimensional cloud-resolving Large Eddy Simulations (LES); each LES simulation represents a different forcing scenario that mimics a range of sites and climates. The observational noise comes from the LES and uses artificial noise inflation to reasonably capture the structural model error. EKI was used to find optimal parameter values. In this experiment, we use an emulator tuned with input-output samples generated by the iterates of the EKI algorithm within CES to approximate the full posterior distribution.
@@
Figure 6: Application of CES to the EDMF problem. Displayed are corner-plots showing slices through the joint distribution of five physically meaningful parameters calibrated to data. Orange (blue) contours: the one, two, and three standard deviation density of the posterior distribution obtained using GP-CES (RF-CES) after hyperparameter tuning.
@@
Figure 6 displays pairwise slices through the five-dimensional posterior distribution for the geophysical problem. Visualized are the final $10^{5}$ steps of MCMC using the likelihood calculated from (53). In orange contours, the GP-CES algorithm is used. In blue contours, the RF-CES algorithm is used. Reasonable qualitative agreement is shown across the support of the approximate posterior obtained from either approach. However, the GP-CES distribution contain some fine-scale artifacts, while the RF-CES distribution appears smoother overall. Such artifacts appear to be due to learning small lengthscales in the GP hyperparameter tuning algorithm rather than physically meaningful features in the posterior distribution. The GP-CES posterior was also slower to converge; four times as many samples of burn-in were taken for GP-CES compared to those taken for RF-CES. Likewise, pairwise correlation between several variables such as entrainment (entrainment_factor) and the stability coefficient (static_stab_coeff) are more clearly seen in the vector-valued RF setting.
Further details of the implementation of this experiment relating to EKI configuration and convergence are provided in Appendix A.5. In that appendix, we also show numerical results using an untuned RF emulator and a tuned RF emulator with a separable feature distribution. One caveat pertaining to the results of the present section is the finding that the structure of the feature distribution plays a significant role in the shape of the posterior distribution (as does the choice of covariance kernel for GP). The design choices (e.g., covariance structure and rank) we selected were based on $L^{2}$ errors of the tuned predictors evaluated on a test set, as reported in Table 6. However, these forward map errors were not as reliable at predicting the skill in sampling the posterior distribution in this experiment, as suggested by the supplementary Figure 12.